Omar is 4 years younger than Christopher. For the last two years, Christopher and Omar have been friends. Three years ago, Christopher was 3 times as old as Omar.
How old is Christopher now?

Answer by JulietG(1812)

C = M + 4 [Omar is 4 years younger than Christopher.]
C-3 = 3(M-3) [Three years ago, Christopher was 3 times as old as Omar.]
Distribute
C-3 = 3M-9
Substitute the known value of C from the first equation into the second.
C-3 = 3M-9
(M+4)-3=3M-9
M+1 = 3M - 9
Subtract M from each side
1 = 2M - 9
Add 9 to each side
10 = 2M
Divide each side by 2
5 = M
Omar is 5. If Omar is 5, then Christopher is 4 years older, or 9.
.
Prove it...
"Three years ago, Christopher was 3 times as old as Omar."
Three years ago, Christopher was 6 (9-3).
Three years ago, Omar was 2 (5-3)
6 is indeed 3 times as many as 2.
.
"For the last two years, Christopher and Omar have been friends." This information is superfluous, and only in there to see if you are smart enough to ignore it.